when you first walk into weaver there are 2 large regular octogonal pillars. the edges are 6.5 and they are 9 feet tall. how much granite was needed to build these pillars??: when you first walk into weaver there are 2 large regular octogonal pillars. the edges are 6.5 and they are 9 feet tall. how much granite was needed to build these pillars??
Answer by nerdybill(1129)

To solve this problem, you need to know the volume of each pillar. The volume is "area of the base" times "height".
.
"area of the base" = "area of an octagon"
.
Without trying to derive the formula of an octagon, I'll just use it. But, for reference, you might want to check out:
http://www.mathreference.com/geo,hex.html
.
"area of an octagon" = 2(1+sqrt(2))S^2
where S=length of one edge
.
So, plugging in your numbers to:
"area of an octagon" = 2(1+sqrt(2))S^2
"area of an octagon" = 2(1+sqrt(2))6.5^2
"area of an octagon" = 2(1+sqrt(2))42.45
"area of an octagon" = 2(1+1.414)42.45
"area of an octagon" = 2(2.414)42.45
"area of an octagon" = (4.828)42.45
"area of an octagon" = 204 sq ft
.
Now, to get the volume we multiply the area times the height:
"vol of one pillar" = "area of an octagon" * "height of pillar"
"vol of one pillar" = 204 * 9
"vol of one pillar" = 1836 cubic feet
.
Since we have 2 pillars, we multiply the above by 2:
2 * 1836 = 3672 cubic feet
.
That's how much granite you need.